OpenStudy (anonymous):

find (g.f)(4) for the functions below f(x) = x+3, g(x) =1/x^2

OpenStudy (anonymous):

\[f(4)=4+3=7\] \[g(7)=\frac{1}{7^2}=\frac{1}{49}\]

OpenStudy (anonymous):

that was \[g\circ f(4)\] if it is \[g\times f(4)\] then the answer is \[g(4)\times f(4)=\frac{1}{49}\times 7=\frac{1}{7}\]